OpenStudy (anonymous):

Use the quadratic formula to solve 2y^2 – 10y + 8 = 0 A.{-4,-1} B.{4, –1} C.{–4, 1} D.{4, 1}

OpenStudy (tkhunny):

Personally, I would divide by 2. \(y^{2} - 5y + 4 = 0\) This gives a = 1, b = -5, and c = 4. Go!

OpenStudy (vishweshshrimali5):

You can also factorise the equation tkhunny obtained.

OpenStudy (tkhunny):

No, there will be no factoring. If the problem statement said "by any method", that would be fine. The requirement is to use the Quadratic Formula. You will see that the Discriminant turns out to be a perfect square. This should tell you that you could have factored it, has we been allowed to do so.

OpenStudy (anonymous):

i solved it the way @tkhunny had said and i got the answer of {5.7,-.7}, I don't know where i went wrong solving it.

OpenStudy (vishweshshrimali5):

Well let me solve it in this way: \(y^2 - 5y + 4 = 0\) \(\implies y^2 - y - 4y + 4 = 0\) \(\implies y(y-1) - 4 (y-1) = 0\) \(\implies (y-4)(y-1) = 0\) \(\implies y = {1, 4}\)

OpenStudy (vishweshshrimali5):

But as @tkhunny said, the question mentions use of quadratic formula so use it only. I only used the above method for verification.

OpenStudy (tkhunny):

Is there some problem with following instructions? This gives a = 1, b = -5, and c = 4 \(b^{2} - 4ac = (-5)^{2} - 4(1)(4) = 25 - 16 = 9 \rightarrow \sqrt{9} = 3\) \(y = \dfrac{5+3}{2} = 8/2 = 4\) \(y = \dfrac{5-3}{2} = 2/2 = 1\) What you should do is write things down. Make your work clear enough to follow. Simply doing it another way, so you are not actually required to do the assignment is NOT a good way to go. Your suggestion that you managed two wrong answers is just not good enough to allow anyone to help you. SHOW YOUR WORK! Trust me on this.

OpenStudy (anonymous):

I get it now, at the beginning I was trying to solve it with -25 insted of just 25. thank you @tkhunny

OpenStudy (tkhunny):

Next time, YOU write clearly so that others can understand. We already know I can do it.